OpenStudy (anonymous):

Solve 2sin2A = 1 for 0º≤A≤360º

OpenStudy (anonymous):

start by dividing by 2. That way you get \[\sin 2A=\frac{ 1 }{ 2 }\] Now, which radians give you sin=1/2?

OpenStudy (anonymous):

2sin^2A = 1, the sin^2 is what im confused by.

OpenStudy (anonymous):

\[\sin ^{2}??\]

OpenStudy (anonymous):

yesum.

OpenStudy (anonymous):

is the original equation squared?

OpenStudy (anonymous):

2sin^2A = 1, that's the original equation.

OpenStudy (anonymous):

oohhh ok. lemme rewrite that

OpenStudy (anonymous):

it would be: \[\sin ^{2}A=\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

ok, how I learned it, solve for A as if it were not squared. Simply, ignore the squared until the end. Now, what gives you what value (in radians? gives you 1/2?

OpenStudy (anonymous):

that doesn't fit into the answers is the only problem

OpenStudy (anonymous):

wait! I just saw my mistake.

OpenStudy (anonymous):